# Reinforce Lifelong Interaction Value of User-Author Pairs for Large-Scale Recommendation Systems

### v1

@@ 4.1.1. Dataset and Evaluation Metrics
The KuaiRand (Gao et al., 2022) is used as the offline experimental dataset to train an offline user simulator. It is a public recommendation dataset containing 27,285 users and 32,038,725 items obtained from Kuaishou app. Therefore, there are hundreds of millions of UA interactions. The trained user simulator is used as the offline environment to mimic the users’ interaction with RS. In details, once simulator receives the recommended item, it returns immediate feedback like click, view and comments, etc. Then it determines whether to send the next request based on a quitting mechanism similar to that in work (Zhang et al., 2024).
We evaluate the performance of compared methods in three aspects, including user satisfaction, author benefits and platform profits:
  • User Satisfaction
@@
4.1.2. Compared Methods
There are five models used as baseline in offline experiment and four variations of our model implemented in ablation experiment:
  • RankingModel: The classic ranking model in RS. Specifically, the DeepFM (Guo et al., 2017) is applied in offline experiments.
  • CQL(Kumar et al., 2020): An offline RL model that introduces a conservative constraint in the Q function update to limit the overly optimistic predictions of actions outside the data distribution.
  • DQN(Van Hasselt et al., 2016): A widely-used RL model that applies DNN as function approximator to estimate the Q-value of each action.
  • TD3(Fujimoto and Gu, 2021): A classic RL model which uses twin critics to reduce the bias, delays the update of policy and smooths the synchrony of target networks.
  • RLUR(Cai et al., 2023): A RL model that aims to optimize the weights of each predicted user feedback when ranking items under the long-term rewards with designed heuristic rewards to overcome the latency and sparsity of long-term rewards.
  • RLIV-UA(w/o MT): The proposed RLIV-UA model without multi-task LIV netowrks.
  • RLIV-UA(w/o MT & SL): The proposed RLIV-UA model without multi-task LIV netowrks and auxiliary supervised learning task.
  • RLIV-UA(w/o AL & SL): The proposed RLIV-UA model without assistance gift label network and auxiliary supervised learning task.
  • RLIV-UA(w/o SL): The proposed RLIV-UA model without auxiliary supervised learning task.
4.1.3. Implementation Details
@@
The assistance gift label network is not applied in offline experiments, because there is no gift signals in KuaiRand dataset.
Besides, the platform profits metrics are only used in online A/B experiments. All models are trained to convergence and their results are the averaged performance of the last 10 epochs.
Moreover, we use the follow LIV and gift LIV in Kwai and use the watch time LIV in Kuaishou and offline experiments.
4.2. Performance Comparison
Table 2. Overall performance of different models in offline recommendation scenario.
Models Session Length Watch Time CTR Diversity
RankingModel 2.0132 59.4812 0.5948 0.0629
CQL 2.2660 58.0753 0.5125 0.1727
DQN 4.0610 49.5454 0.2247 0.7477
TD3 4.7820 52.7157 0.2228 0.8500
RLUR 6.6810 151.0550 0.4519 0.7206
RLIV-UA 12.8860 377.4867 0.5015 0.8827
Table 3. Overall performance of variations of the proposed RLIV-UA model in offline recommendation scenario.
Models Session Length Watch Time CTR Diversity
RLIV-UA(w/o MT & SL) 7.2940 188.0499 0.5229 0.7994
RLIV-UA(w/o MT) 9.2800 248.1804 0.5340 0.8746
RLIV-UA 12.8860 377.4867 0.5015 0.8827
4.3. Ablation Study
The overall performance of RLIV-UA variations in offline recommendation scenario are shown in Table 3. Firstly, compared with the RLIV-UA(w/o MT & SL) variation, the RLIV-UA(w/o MT) achieves relatively high improvement in session length, watch time and diversity, which reflects that the auxiliary supervised learning task can help model learn the LIV of UA pairs more accurately.
4.4. Online A/B Experiments
Firstly, the proposed RLIV-UA model and its variants are deployed on Kwai live-stream feed with over 100 million users and 1 million authors to improve platform revenue from July to October 2024. Through online A/B experiments, the RLIV-UA(w/o AL & SL), RLIV-UA(w/o SL) and RLIV-UA models are successively evaluated under the platform profits metrics compared to RankingModel as baseline. As shown in Table 4, RLIV-UA variations successively perform better in both revenue metrics and new fans metric, which illustrates the practical effectiveness of all parts of RLIV-UA model.
Table 4. Results of the revenue experiment on Kwai live-stream feed.
Models UA Count Weekly Gifted Users Total Revenue New Fans
RankingModel - - - -
RLIV-UA(w/o AL & SL) +2.002% +2.245% +11.903% +2.743%
RLIV-UA(w/o SL) +3.579% +3.417% +23.8065% +6.104%
RLIV-UA +5.849% +5.370% +38.6445% +8.282%
Table 5. Results of the watch time experiment on Kuaishou short-video feed.
Models Watch Time App Usage Time Weekly Retention
RankingModel - - -
RLIV-UA +0.131% +0.083% +0.021%
Figure 6. The output lifelong gift values of different grades of gift amount among UA pairs.
Moreover, the output lifelong gift values corresponding to different grades of gift amount among UA pairs are shown in Fig. 6, illustrating the high positive correlation between the output lifelong gift value and actual revenue.
The proposed RLIV-UA model is also deployed on Kuaishou short-video feed to improve long-term user engagement from November to December 2024. The results are listed in Table 5, which shows that the RLIV-UA model can improve the app usage time by 0.083% and the weekly retention by 0.021%. It should be emphasized that a 0.02% improvement holds statistical significance in the system. It is proved that the RLIV-UA model can not only improve platform revenue but also improve users’ long-term retention.
5. Related work
5.1. RL-Based Recommendation Systems
(Sunehag et al., 2015) is the earliest work that tries to alternate multitask learning ranking model with RL model using DQN to learn the value of all items in the recommended list. Similarly, Chen et al. (Chen et al., 2019) employ a policy-gradient approach in RS and Zhao et al. (Zhao et al., 2018) develop an actor-critic approach for recommending a page of items. However, they are not applied in a real-world recommendation environment with large amount of users and items. Then, more research (Gauci et al., 2018; Zheng et al., 2018) aims to apply the RL model in reality as a substitute with simple network structure. In order to handle the huge number of candidate items, SlateQ (Ie et al., 2019) is proposed to decompose the value of item list into the sum of value of each item under some assumptions. Recently, some literatures (Deffayet et al., 2023; Liu et al., 2023a) use contrastive learning to overcome the curse of dimensionality whose model structures are relatively more complex.
5.2. Long-Term User Engagement in Recommendation Systems
In order to consider the long-term user engagement rather than user’s immediate feedback, some research has increasingly focused on the sequential patterns of user behavior by employing temporal models, such as hidden Markov models and recurrent neural networks (Rendle et al., 2010; Campos et al., 2014; He and McAuley, 2016; Sahoo et al., 2012; Tan et al., 2016; Wu et al., 2017). Besides, some research (Gauci et al., 2018; Shani et al., 2005; Taghipour et al., 2007) use RL to make a long-term planning. However, all the methods are too complex to be applied in practice. Zou et al. (Zou et al., 2019) propose a hierarchical LSTM based Q network to model the complex user behavior and design an S-network to simulate the environment avoiding the instability. Chen et al. (Chen et al., 2021) inspired by exploration research (Mnih, 2016; Bellemare et al., 2016) in RL use a series of exploration methods to improve user experience. Wang et al. (Wang et al., 2022) carefully design the reward function through data analysis to connect the long-term rewards with immediate feedback. While Xue et al. (Xue et al., 2023) propose a framework for learning preferences from user historical behavior sequences, specifically using preferences to automatically train a reward function in an end-to-end manner. Considering all the above methods’ action is to select an item list which may be not practical when the number of item and user is large, Cai et al. (Cai et al., 2023) aim to optimize the weights of each predicted user feedback when ranking items under the long-term rewards with designed heuristic rewards to overcome the latency and sparsity of long-term rewards.
6. Conclusion
7. GenAI Usage Disclosure
We guarantee that no GenAI tools were used in any stage of the research, nor in the writing.

### v2

@@ 4.1.1. Dataset and Evaluation Metrics
We adopt the KuaiRand dataset (Gao et al., 2022) as the foundation for our offline experiments. This public dataset, collected from the Kuaishou app, contains interactions from 27,285 users across 32,038,725 items, resulting in hundreds of millions of UA interaction records. We train an offline user simulator on this dataset to serve as a controllable environment that mimics real user behavior. Specifically, upon receiving a recommended item, the simulator generates immediate feedback signals, such as clicks, watch duration, and comments, and subsequently decides whether to issue the next request based on a probabilistic quitting mechanism, similar to the approach described in (Zhang et al., 2024).
To comprehensively evaluate model performance, we assess three key dimensions: user satisfaction, author benefits, and platform profitability. The detailed illustration of all evaluation metrics is available in Appendix B.1.
  • User Satisfaction is measured by: Session Length (number of requests per session), Watch Time (total viewing duration per session), and CTR (click-through rate).
  • Author Benefits are quantified via: Diversity (variety of recommended authors/content) and New Fans (number of new followers acquired by authors).
  • Platform Profits are evaluated using: UA Count (number of UA pairs exhibiting deep relationships), Weekly Gifted Users (number of users sending virtual gifts per week), Total Revenue, App Usage Time, and Weekly Retention (proportion of users returning weekly).
4.1.2. Compared Methods
We compare our method against a suite of representative baselines, including the classic RankingModel and four established reinforcement learning approaches: CQL (Kumar et al., 2020), DQN (Van Hasselt et al., 2016), TD3 (Fujimoto and Gu, 2021), TD3-UA (Using TD3 to model the SCRI-MDP and the action is a continuous weight value), FeedRec (Zou et al., 2019) and RLUR (Cai et al., 2023). To further validate the contribution of each component in RLIV-UA, we also evaluate four ablated variants: RLIV-UA(w/o MT) (without Multi-Task learning), RLIV-UA(w/o MT & SL) (without Multi-Task and Supervised Learning), RLIV-UA(w/o AL & SL) (without Auxiliary Gift Label Network and Supervised Learning), and RLIV-UA(w/o SL) (without Supervised Learning). The implementation details is available in Appendix B.2.
Note that the assistance gift label network is not applied in offline experiments, because there is no gift signals in KuaiRand dataset. The platform profits metrics are only used in online A/B experiments. All models are trained to convergence, and their results are the averaged performance of the last 10 epochs. Moreover, we use the follow LIV and gift LIV in Kwai and use the watch time LIV in Kuaishou and offline experiments.
4.2. Performance Comparison
4.3. Ablation Study
The overall performance of RLIV-UA variations in offline recommendation scenario are shown in Table 1.
Firstly, compared with the RLIV-UA(w/o MT & SL) variation, the RLIV-UA(w/o MT) achieves relatively high improvement in session length, watch time and diversity, which reflects that the auxiliary supervised learning task can help model learn the LIV of UA pairs more accurately.
As shown in Figure 7, the RLIV-UA model with $\gamma=0.9$ achieves the best performance in session length and watch time and all the models achieve better performance compared with baselines, which shows the RLIV-UA model is not sensitive to the parameter $\gamma$.
(a) The session length of RLIV-UA with different $\gamma$.
(b) The watch time of RLIV-UA with different $\gamma$.
Figure 7. The parameter sensitivity experiment of $\gamma$.
Table 2. Results of the revenue experiment on Kwai live-stream feed.
Models
UA Count
Weekly Gifted Users
Total Revenue
New Fans
RankingModel - - - -
RLIV-UA(w/o AL & SL) +2.00% CI:[0.97%, 5.08%] +2.25% CI:[0.46%, 4.08%] +11.90% CI:[5.48%, 32.32%] +2.74% CI:[0.60%, 4.92%]
RLIV-UA(w/o SL) +3.58% CI:[1.18%, 5.36%] +3.42% CI:[0.23%, 5.11%] +23.81% CI:[15.93%, 37.52%] +6.10% CI:[0.10%, 10.15%]
RLIV-UA(w/o ASA) +2.02% CI:[1.71%, 5.74%] +1.68% CI:[0.26%, 3.10%] +15.21% CI:[0.36%, 21.34%] +5.22% CI:[0.36%, 10.08%]
RLIV-UA +5.85% CI:[1.34%, 8.93%] +5.37% CI:[0.11%, 8.71%] +38.64% CI:[13.43%, 60.31%] +8.28 CI:[0.91%, 14.38%]
4.4. Online A/B Experiments
The proposed RLIV-UA model and compared methods were deployed on the Kwai live-streaming feed, a platform serving over 100 million users and 1 million content authors, to optimize platform revenue from July to October 2024. For rigorous evaluation, we randomly assigned 20% of users (over 20 million) to the treatment group, with the remaining 80% serving as the control. All reported confidence intervals (CIs) in Table 2 are 95% two-sided CIs, computed based on the Central Limit Theorem, ensuring statistical robustness given the massive sample size.
As shown in Table 2, the performance of RLIV-UA and its ablated variants improves progressively: RLIV-UA(w/o AL&SL) → RLIV-UA(w/o SL) → full RLIV-UA. This consistent uplift across key business metrics, including Total Revenue +38.64%, Weekly Gifted Users +5.37%, UA Count +5.85%, and New Fans +8.28%, demonstrates the practical effectiveness and additive value of each component in our framework. Notably, we also evaluate a variant, RLIV-UA(w/o ASA), which forgoes the ASA module and instead constructs RL training samples using only stitched historical interactions. This variant underperforms the full RLIV-UA model across all metrics, highlighting ASA’s critical role in mitigating sample sparsity and bias. Without ASA, the model cannot learn from first-time or infrequent UA interactions, severely limiting its ability to optimize lifelong value. This confirms ASA is not optional, but essential for industrial-scale deployment.
To further validate generalizability, we deployed RLIV-UA on the Kuaishou short-video feed from November to December 2024, again using a 20% user cohort for treatment. Results show statistically significant improvements in long-term user engagement: +0.131% in average watch time, +0.083% in daily app usage duration, and +0.021% in weekly retention rate. Notably, in a system of Kuaishou’s scale, even a 0.02% gain in retention is considered highly significant, reflecting millions of additional active users. These results confirm that RLIV-UA not only drives immediate revenue growth but also fosters sustainable, long-term platform health by strengthening user-author relationships.
5. Related work
5.1. RL-Based Recommendation Systems
(Sunehag et al., 2015) is the earliest work that tries to alternate multitask learning ranking model with RL model using DQN to learn the value of all items in the recommended list. Similarly, Chen et al. (Chen et al., 2019) employ a policy-gradient approach in RS and Zhao et al. (Zhao et al., 2018) develop an actor-critic approach for recommending a page of items. However, they are not applied in a real-world recommendation environment with large amount of users and items. Then, more research (Gauci et al., 2018; Zheng et al., 2018) aims to apply the RL model in reality as a substitute with simple network structure. In order to handle the huge number of candidate items, SlateQ (Ie et al., 2019) is proposed to decompose the value of item list into the sum of value of each item under some assumptions. Other literatures (Deffayet et al., 2023; Liu et al., 2023a) use contrastive learning to overcome the curse of dimensionality whose model structures are more complex.
5.2. Long-Term User Engagement in Recommendation Systems
In order to consider the long-term user engagement rather than user’s immediate feedback, some research has increasingly focused on the sequential patterns of user behavior by employing temporal models, such as hidden Markov models and recurrent neural networks (Rendle et al., 2010; Campos et al., 2014; He and McAuley, 2016; Sahoo et al., 2012; Tan et al., 2016; Wu et al., 2017). Besides, some research (Gauci et al., 2018; Shani et al., 2005; Taghipour et al., 2007) use RL to make a long-term planning. However, all the methods are too complex to be applied in practice. (Zou et al., 2019) propose a hierarchical LSTM based Q network to model the complex user behavior and design an S-network to simulate the environment avoiding the instability. (Chen et al., 2021) inspired by exploration research (Mnih, 2016; Bellemare et al., 2016) in RL use a series of exploration methods to improve user experience. (Wang et al., 2022) carefully design the reward function through data analysis to connect the long-term rewards with immediate feedback. While (Xue et al., 2023) propose a framework for learning preferences from user historical behavior sequences, specifically using preferences to automatically train a reward function in an end-to-end manner. Considering all the above methods’ action is to select an item list which may be not practical when the number of item and user is large, (Cai et al., 2023) aim to optimize the weights of each predicted user feedback when ranking items under the long-term rewards with designed heuristic rewards to overcome the latency and sparsity of long-term rewards.
6. Conclusion
As future work, we plan to explore the application of RLIV-UA’s core components, particularly the SCRI-MDP formulation, ASA method, and MTCL architecture , in other domains such as e-commerce (user-merchant) and music streaming (user-artist), where optimizing sparse, long-term, bidirectional value is equally critical. This will require careful adaptation of reward functions and interaction definitions to fit each domain’s unique dynamics.
Appendix A Proofs
A.1. Equivalence to augmented MDP
Theorem A.1 (Equivalence to augmented MDP).
Consider a Sparse Cross-Request Interaction MDP $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\Delta\rangle$ with interaction gap variable $\Delta^{t}_{\mathrm{ua}}\geq 1$. Define an augmented state $\bar{s}_{t}=(s^{t}_{\mathrm{ua}},\Delta^{t}_{\mathrm{ua}})$ where $\Delta^{t}_{\mathrm{ua}}$ is treated as part of the state vector, forming an augmented state space $\bar{\mathcal{S}}=\mathcal{S}\times\mathbb{N}$. Then $\mathcal{M}$ is equivalent to a standard MDP $\bar{\mathcal{M}}=\langle\bar{\mathcal{S}},\mathcal{A},\bar{\mathcal{P}},\mathcal{R},\gamma\rangle$, where the transition kernel is defined as:
(16) $\bar{P}((\mathbf{s}^{\prime},\Delta^{\prime})\mid(\mathbf{s},\Delta),a)=P(\mathbf{s}^{\prime}_{\mathrm{ua}},\Delta^{\prime}_{\mathrm{ua}}\mid\mathbf{s}_{\mathrm{ua}},\Delta_{\mathrm{ua}},a_{\mathrm{ua}}),$
and the reward function $\mathcal{R}$ remains identical. Any optimal policy $\pi^{*}$ in $\mathcal{M}$ corresponds to an optimal policy in $\bar{\mathcal{M}}$ under this mapping.
Proof.
We aim to prove that an optimal policy in the SCRI-MDP $\mathcal{M}$ corresponds to an optimal policy in the augmented MDP $\bar{\mathcal{M}}$. We proceed via proof by contradiction.
Let $\pi^{*}:\mathcal{S}\times\mathbb{N}\to\mathcal{A}$ be an optimal deterministic policy for the SCRI-MDP $\mathcal{M}$. By definition, for all $(s,\Delta)\in\mathcal{S}\times\mathbb{N}$ and for any other policy $\pi$, the value function satisfies:
(17) $V^{\pi^{*}}(s,\Delta)\geq V^{\pi}(s,\Delta)$
Now, consider the corresponding policy $\bar{\pi}^{*}$ in the augmented MDP $\bar{\mathcal{M}}$, defined such that $\bar{\pi}^{*}(a|\bar{s})=\pi^{*}(a|s,\Delta)$ for $\bar{s}=(s,\Delta)$. Assume, for the sake of contradiction, that $\bar{\pi}^{*}$ is not optimal in $\bar{\mathcal{M}}$. This implies that there exists another policy $\bar{\pi}^{\prime}:\bar{\mathcal{S}}\to\mathcal{A}$ and some augmented state $\bar{s}_{0}=(s_{0},\Delta_{0})$ such that:
(18) $V^{\bar{\pi}^{\prime}}(s_{0},\Delta_{0})>V^{\bar{\pi}^{*}}(s_{0},\Delta_{0})$
Define a policy $\pi^{\prime}$ for the original SCRI-MDP $\mathcal{M}$ by $\pi^{\prime}(a|s,\Delta):=\bar{\pi}^{\prime}(a\mid(s,\Delta))$. Since both MDPs share the same reward function $\mathcal{R}$, discount factor $\gamma$, and the dynamics of $\mathcal{M}$ are fully encoded in $\bar{\mathcal{P}}$ through the inclusion of $\Delta$ in the state, the expected return from any initial state-action pair is identical under corresponding policies. Specifically, the Bellman equations for both frameworks satisfy:
For the SCRI-MDP:
(19) $Q^{\pi}(s,\Delta,a)$ $=\mathcal{R}(s,\Delta,a)+$
$\gamma\sum_{s^{\prime},\Delta^{\prime}}P(s^{\prime},\Delta^{\prime}\mid s,\Delta,a)\sum_{a^{\prime}}\pi(a^{\prime}\mid s^{\prime},\Delta^{\prime})Q^{\pi}(s^{\prime},\Delta^{\prime},\pi)$
For the augmented MDP:
(20) $Q^{\bar{\pi}}((s,\Delta),a)=\mathcal{R}(s,\Delta,a)+$
$\gamma\sum_{s^{\prime},\Delta^{\prime}}\bar{P}((s^{\prime},\Delta^{\prime})\mid(s,\Delta),a)\sum_{a^{\prime}}\bar{\pi}(a^{\prime}\mid s^{\prime},\Delta^{\prime})Q^{\bar{\pi}}((s^{\prime},\Delta^{\prime}),\bar{\pi}))$
Since $\bar{P}((s^{\prime},\Delta^{\prime})\mid(s,\Delta),a)=P(s^{\prime},\Delta^{\prime}\mid s,\Delta,a)$ by construction, and the reward function is identical, it follows that:
(21) $V^{\pi^{\prime}}(s,\Delta)=V^{\bar{\pi}^{\prime}}(s,\Delta)\quad\text{and}\quad V^{\pi^{*}}(s,\Delta)=V^{\bar{\pi}^{*}}(s,\Delta)$
for all $(s,\Delta)\in\mathcal{S}\times\mathbb{N}$.
Substituting these identities into our earlier inequality yields:
(22) $V^{\pi^{\prime}}(s_{0},\Delta_{0})>V^{\pi^{*}}(s_{0},\Delta_{0})$
This contradicts the assumption that $\pi^{*}$ is optimal for $\mathcal{M}$. Therefore, the assumption that $\bar{\pi}^{*}$ is not optimal in $\bar{\mathcal{M}}$ must be false. Hence, $\bar{\pi}^{*}$ is indeed an optimal policy for $\bar{\mathcal{M}}$.
Conversely, let $\bar{\pi}^{*}$ be an optimal deterministic policy for the augmented MDP $\bar{\mathcal{M}}$. Then, we have:
(23) $V^{\bar{\pi}^{*}}(\bar{s})\geq V^{\bar{\pi}}(\bar{s})$
for all $\bar{s}\in\bar{\mathcal{S}}$ and for any other policy $\bar{\pi}$.
Define a policy $\pi^{*}$ for the SCRI-MDP $\mathcal{M}$ by $\pi^{*}(a|s,\Delta):=\bar{\pi}^{*}(a\mid(s,\Delta))$. Assume, for contradiction, that $\pi^{*}$ is not optimal in $\mathcal{M}$. Then there exists another policy $\pi^{\prime}$ and some state $(s_{0},\Delta_{0})$ such that:
(24) $V^{\pi^{\prime}}(s_{0},\Delta_{0})>V^{\pi^{*}}(s_{0},\Delta_{0})$
Define $\bar{\pi}^{\prime}$ for $\bar{\mathcal{M}}$ by $\bar{\pi}^{\prime}(a\mid(s,\Delta)):=\pi^{\prime}(a\mid s,\Delta)$. Using the same argument as before, we have:
(25) $V^{\bar{\pi}^{\prime}}(s_{0},\Delta_{0})>V^{\bar{\pi}^{*}}(s_{0},\Delta_{0})$
which contradicts the optimality of $\bar{\pi}^{*}$ in $\bar{\mathcal{M}}$. Therefore, $\pi^{*}$ must be optimal in $\mathcal{M}$.
This establishes a one-to-one correspondence between optimal policies in $\mathcal{M}$ and $\bar{\mathcal{M}}$. Consequently, the SCRI-MDP $\mathcal{M}$ is equivalent to the standard MDP $\bar{\mathcal{M}}$ under the augmented state formulation, completing the proof. ∎
A.2. Error Bound for ASA
Theorem A.2 (Error Bound for Approximate State Transitions).
Let $Q^{*}$ be the optimal Q-function of the SCRI-MDP $\mathcal{M}$, and let $\hat{Q}$ be the Q-function learned using the ASA with maximum prediction error $\epsilon=\max_{s,a,r}\left|\hat{s}^{t+1}-s^{t+1}\right|$. If the ASA error $\epsilon$ is bounded and the reward function is Lipschitz continuous with constant $L_{r}$, then the difference between $\hat{Q}$ and $Q^{*}$ is bounded by:
(26) $\left|\hat{Q}(s,a)-Q^{*}(s,a)\right|\leq\frac{L_{r}\epsilon}{(1-\gamma)^{2}}$
for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, where $\gamma\in[0,1)$ is the discount factor.
Proof.
Let’s denote the Bellman operator for the true MDP as $\mathcal{T}$, and for the approximate MDP with ASA as $\hat{\mathcal{T}}$. The Bellman operators are defined as:
(27) $\mathcal{T}Q(s,a)$ $=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\left[\max_{a^{\prime}}Q(s^{\prime},a^{\prime})\right]$
$\hat{\mathcal{T}}Q(s,a)$ $=r(s,a)+\gamma\mathbb{E}_{\hat{s}^{\prime}\sim\hat{P}(\cdot|s,a)}\left[\max_{a^{\prime}}Q(\hat{s}^{\prime},a^{\prime})\right]$
where $P$ is the true transition probability and $\hat{P}$ is the approximate transition probability induced by ASA.
The fixed points of these operators are the optimal Q-function $Q^{*}$ and the approximate Q-function $\hat{Q}$, respectively:
(28) $Q^{*}$ $=\mathcal{T}Q^{*}$
$\hat{Q}$ $=\hat{\mathcal{T}}\hat{Q}$
We want to bound $\left|\hat{Q}(s,a)-Q^{*}(s,a)\right|$. Consider:
(29) $\left|\hat{Q}(s,a)-Q^{*}(s,a)\right|=\left|\hat{\mathcal{T}}\hat{Q}(s,a)-\mathcal{T}Q^{*}(s,a)\right|$
$\leq\left|\hat{\mathcal{T}}\hat{Q}(s,a)-\hat{\mathcal{T}}Q^{*}(s,a)\right|+\left|\hat{\mathcal{T}}Q^{*}(s,a)-\mathcal{T}Q^{*}(s,a)\right|$
The first term can be bounded using the contraction property of the Bellman operator:
(30) $\left|\hat{\mathcal{T}}\hat{Q}(s,a)-\hat{\mathcal{T}}Q^{*}(s,a)\right|$ $\leq\gamma\left|\hat{Q}-Q^{*}\right|$
For the second term, we have:
(31) $\left|\hat{\mathcal{T}}Q^{*}(s,a)-\mathcal{T}Q^{*}(s,a)\right|$
$=\gamma\left|\mathbb{E}_{\hat{s}^{\prime}\sim\hat{P}(\cdot|s,a)}\left[\max_{a^{\prime}}Q^{*}(\hat{s}^{\prime},a^{\prime})\right]-\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\left[\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime})\right]\right|$
Since $Q^{*}$ is Lipschitz continuous with constant $L_{Q}=\frac{L_{r}}{1-\gamma}$ (a standard result in RL), and the ASA error is bounded by $\epsilon$, we have:
(32) $\left|\hat{\mathcal{T}}Q^{*}(s,a)-\mathcal{T}Q^{*}(s,a)\right|$ $\leq\gamma L_{Q}\epsilon=\gamma\frac{L_{r}}{1-\gamma}\epsilon$
Combining these results:
(33) $\left|\hat{Q}(s,a)-Q^{*}(s,a)\right|$ $\leq\gamma\left|\hat{Q}-Q^{*}\right|+\gamma\frac{L_{r}}{1-\gamma}\epsilon$
Taking the supremum over all $(s,a)$:
(34) $\left|\hat{Q}-Q^{*}\right|$ $\leq\gamma\left|\hat{Q}-Q^{*}\right|+\gamma\frac{L_{r}}{1-\gamma}\epsilon$
$(1-\gamma)\left|\hat{Q}-Q^{*}\right|$ $\leq\gamma\frac{L_{r}}{1-\gamma}\epsilon$
$\left|\hat{Q}-Q^{*}\right|$ $\leq\frac{L_{r}\epsilon}{(1-\gamma)^{2}}$
This completes the proof. ∎
Appendix B Experimental Settings and Algorithm Pseudocode
B.1. Evaluation Metrics
The detailed evaluation metrics of compared methods are shown as follows:
@@ B.1. Evaluation Metrics
B.2. Implementation Details
@@
B.3. Algorithm Pseudocode
As shown in Algorithm 1, the training process of the RLIV-UA is divided into two steps: The pre-training of the state predictor $f_{\theta}$ and the joint training of the LIV networks and the tate predictor.
Algorithm 1 The RLIV-UA Algorithm
1:Input: Initial parameters of all the Q networks $\phi_{i},i=1,...,n$, initial parameters of state predictor $\theta$, learning rate $\lambda$, batch size for updating $B$ and the update rate of target networks $\tau$
2:Initialize the target networks with $\phi^{\prime}_{i}\leftarrow\phi_{i}$
3:Initialize the replay memory buffer of Q networks $D\leftarrow\emptyset$
4:Repeat:
5:Step1: The pre-training of state predictor
6:for $\text{epoch}=1$ to train_epochs do
7:  for each interact_step $t$ of each user-author pair $\mathrm{ua}$ do
8:   Observe the current state $\mathbf{s}^{t}_{\mathrm{ua}}$.
9:   Retrieve the last interaction from user historical behaviors sequences $\mathbf{s}^{t-1}_{\mathrm{ua}},a^{t-1}_{\mathrm{ua}},r^{t-1}_{\mathrm{ua}}$.
10:   Conduct gradient update for $f_{\theta}$ $\theta\leftarrow\theta-\lambda\nabla_{\theta}\mathcal{L}_{\theta}(\theta)$
11:  end for
12:end for
13:Step2: The join-training of LIV networks
14:for $\text{epoch}=1$ to train_epochs do
15:  for each interact_step $t$ of each user-author pair $\mathrm{ua}$ do
16:   Observe the current state $\mathbf{s}^{t}_{\mathrm{ua}}$, action $a^{t}_{\mathrm{ua}}$ and rewards $r^{t}_{\mathrm{ua},i},i=1,...,n$.
17:   Input them into the state predictor $f_{\theta}$ to predict the next state $\hat{\mathbf{s}}^{t+1}_{\mathrm{ua}}$.
18:   $D$ $\leftarrow$ $D\cup(\mathbf{s}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}},r^{t}_{\mathrm{ua}},\hat{\mathbf{s}}^{t+1}_{\mathrm{ua}})$
19:   Retrieve the last interaction from user historical behaviors sequences $\mathbf{s}^{t-1}_{\mathrm{ua}},a^{t-1}_{\mathrm{ua}},r^{t-1}_{\mathrm{ua}}$.
20:   Conduct gradient update for $f_{\theta}$ $\theta\leftarrow\theta-\lambda\nabla_{\theta}\mathcal{L}_{\theta}(\theta)$
21:   if the number of samples in $D$ reaches $B$ then
22:     sample a batch from $D$ and calculate total loss $\mathcal{L}$ shown in 15
23:     $\phi_{i}\leftarrow\phi_{i}-\lambda\nabla_{\phi_{i}}\mathcal{L}(\phi_{i})$
24:     $\phi^{\prime}_{i}\leftarrow(1-\tau)\phi_{i}+\tau\phi^{\prime}_{i}$
25:   end if
26:  end for
27:end for
28:Return The final Q networks parameters
Appendix C Data Analysis
Firstly, we observe that the lifelong UA interactions have strong connection with the ultimate platform revenue. As shown in Fig. 8(a), the count of UA pairs with both follow and frequent gift (”deep”) relationship is positively correlated with the total revenue value. On the other hand, as users continue to interact with the authors with ”deep” relationship, its conversion rate will increase, as shown in Fig. 8(b). Hence, it provides insights to improve the platform profits by modeling the LIV of UA pairs.
(b) The conversion rate of UA pairs with ”deep” relationship.
Figure 9. The output lifelong revenue values of some user-author pairs as their relationship becomes deeper.
In particular, to more clearly evaluate how the RLIV-UA model optimizes the LIV of UA pairs a specific example is shown in Fig. 9. We collect the lifelong revenue value (Q value) output by the RLIV-UA model with progressive UA interactions from click, watch to interact, and gift. It shows that as the UA output lifelong revenue value of user-author pair is progressively increasing as their relationship becomes deeper.
Appendix D More Discussion of ASA
D.1. The rule-based approximation
The details of the rule-based approximation are as follows.
Specifically, the dynamic features $\mathbf{f}_{\mathrm{ua}}^{t}$ in the current state $\mathbf{s}^{t}_{\mathrm{ua}}$ is defined as the counts of several kinds of interactions between $\mathrm{ua}$. Note that $\mathbf{f}_{\mathrm{ua}}^{t+1}$ can be approximated based on user’s current rewards $r^{t}_{\mathrm{ua}}$ and current dynamic UA features $\mathbf{f}_{\mathrm{ua}}^{t}$:
[…]
Since the SCRI-MDP only focuses on the interactions and state transitions between UA pair like $\mathrm{ua}$, it ignores interactions between user $u$ and items of other authors. Under the above assumption of the SCRI-MDP, the next state $\hat{\mathbf{s}}^{t+1}_{ua}$ can be derived by the dynamic features $\mathbf{f}^{t+1}_{ua}$.
D.2. Offline evaluation of ASA
Table 4. The evaluation loss of the adjacent state approximation compared with rule-based approximation.
Methods MSE loss Cosine Similarity loss
rule-based approximation 0.718 0.743
proposed ASA method 0.502 0.630
To demonstrate the effectiveness of the ASA method, we conduct a comparative evaluation against a rule-based baseline on an offline test set. The full dataset is randomly split into training and test sets in an 8:2 ratio. As shown in Table 4, the ASA method achieves significantly lower prediction error across all key state features, confirming its superior ability to model the complex, long-term dynamics of user-author interactions compared to the heuristic rule-based approach.